### v1

@@ 3.2 Escaping Saddle Points
Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function. For any strict saddle point ${\mathbf{x}}\in\mathrm{Crit}(f)$ with value $f({\mathbf{x}})=c$ there exists $N>0$ and a neighborhood $U$ of ${\mathbf{x}}$ such that
$\Pr(f({\mathbf{x}}_{i+N})<c\,|\,{\mathbf{x}}_{i}\in U)>\delta.$
@@
This follows immediately from Corollary 3.7, Lemma 3.8 and Lemma 3.9. ∎
@@
Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function. Further assume that $f$ has compact sublevel sets and that $\mathbf{x}_{0}\in M$ is not a critical point of $f$. Then the probability that $\mathbf{x}_{i}$ converges to a critical submanifold of a strict saddle points as $i\to\infty$ is zero.
Proof.
Consider some critical value $c$ of $f$ and let $C\subset M$ be a connected component of the corresponding critical submanifold made up of strict saddle points, cf. [48, Def. 2.41]. Note that $C$ is compact. We will show that the probability that $\mathbf{x}_{i}$ converges to $C$ is zero. Indeed let ${\mathbf{z}}\in C$ be arbitrary. Then, by Lemma 3.10, there exists an open set $U\subset M$ containing ${\mathbf{z}}$ such that $\Pr(f(\mathbf{x}_{i+N})<c\,|\,\mathbf{x}_{i}\in U)>\delta$ for some $\delta>0$ and $N>0$.
Moreover, by compactness we can find a finite collection of such open sets covering $C$. Hence it suffices to show that the probability that $\mathbf{x}_{i}$ has an accumulation point in $C\cap U$ is zero. Consider a sequence $\mathbf{x}_{i}$ and let $\mathbf{z}\in C\cap U$ be an accumulation point. Consider two balls $B_{R_{1}}(\mathbf{z}),B_{R_{2}}(\mathbf{z})$ centered on $\mathbf{z}$ with $0<R_{2}<R_{1}$ and both balls contained in $U$. Choosing $R_{1}$ and $R_{2}/R_{1}$ small enough we can guarantee that $\mathbf{x}_{0}\notin B_{R_{1}}(\mathbf{z})$ and that any realization needs at least $N$ steps to traverse the spherical shell $B_{R_{1}}\setminus B_{R_{2}}$. Now choosing a sequence of strictly decreasing $R_{k}$ with $R_{1}$ and all $R_{2k}/R_{2k-1}$ small enough as above we see that the sequence must traverse all shells $B_{R_{2k-1}}\setminus B_{R_{2k}}$
This shows that any sequence with accumulation point in $C\cap U$ must, infinitely often, spend $N$ consecutive steps in $U$. By Lemma 3.10 the probability that this happens is zero. Here we used the fact that by Markovianity the behaviour of each sequence of $N$ consecutive steps behaves independently. This concludes the proof. ∎
3.3 Almost Sure Convergence
The results proven so far guarantee that the algorithm converges almost surely to the set of local minima:
Proposition 3.12.
Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with compact sublevel sets. Assume that $\mathbf{x}_{0}$ is not a critical point. Then for stepsize $\eta\leq 1/\ell$ the randomized gradient descent algorithm converges almost surely to the set of local minima.
Proof.
Since by Lemma 3.4 the function value cannot increase, and since by compactness of sublevel sets the function is lower bounded, the function value must (surely) converge to some value $f^{\star}$. By Corollary 3.5, $f^{\star}$ is almost surely a critical value. By Lemma 3.11 the algorithm almost surely does not converge towards a strict saddle point. Hence we converge almost surely to the set of local minima. ∎
In order to prove convergence to a single local minimum, we first need the following lemma.
Lemma 3.13.
Let $\mathcal{E}_{i}$ be probability distributions with values in the interval $[0,1]$ and let $E_{i}\sim\mathcal{E}_{i}$ for $i\in\mathbb{N}$. Let $\varepsilon>0$. We assume that there is $q>0$ such that
$\Pr(E_{i}\geq\varepsilon\,|\,E_{1}=e_{1},\ldots,E_{i-1}=e_{i-1})\geq q$
for all $i\in\mathbb{N}$ and for all $e_{1},\ldots,e_{i-1}\in[0,1]$. Then for any $\alpha\in(0,q\ln\tfrac{1}{1-\varepsilon})$ it holds that almost surely
@@ 3.3 Almost Sure Convergence
First recall the following basic fact. Consider a sequence of i.i.d. biased coin tosses $X_{i}\in\{0,1\}$ with $\Pr(X_{i}=1)=q$. By the strong law of large numbers [30, Thm. 20.1] the average $\overline{X}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}X_{i}$ converges almost surely to the expectation $q$. Hence for any $p\in(0,q)$, there almost surely exists some $n_{0}$ large enough such that $\overline{X}_{n}\geq p$ for all $n>n_{0}$.
Let any value $n$, and index set $I=\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,n\}$ (ordered increasingly) be given. Then
@@
Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with $f$ has compact sublevel sets. Further assume that ${\mathbf{x}}_{0}$ is not a critical point of $f$. Then, for stepsize $\eta\leq 1/\ell$, the randomly projected gradient descent algorithm converges almost surely to a local minimum.
@@
The idea is the proof is as follows: first we show that almost surely the distance to the set of local minima decreases exponentially. This then implies that the size of each step decreases exponentially as well, and hence the algorithm converges absolutely to a local minimum.
Proposition 3.12 shows that ${\mathbf{x}}_{i}$ converges almost surely to some connected set of local minima denoted $C$. Let ${\mathbf{z}}\in C$ be such a local minimum, and let $H({\mathbf{z}})$ denote the Hessian at ${\mathbf{z}}$, and let $a_{\min}({\mathbf{z}})$ denote the smallest non-zero eigenvalue of $H({\mathbf{z}})$. By compactness and continuity $a_{\min}({\mathbf{z}})$ has a minimum on $C$, which we denote $a_{\min}$. In a small enough neighborhood of ${\mathbf{z}}$, Lemma 3.9 shows that
@@
Together with Lemma 3.4 we find
$f({\mathbf{x}}_{i+1})\leq\Big{(}1-((\mathbf{u}_{i})_{N})^{2}\delta 2a_{\min}(\mathbf{z})\Big{)}f({\mathbf{x}}_{i})$
where $\delta=\eta(1-\tfrac{\ell\eta}{2})$.
By Lemma 3.13 this shows that there is some $\alpha>0$ such that $f({\mathbf{x}}_{i})\leq f({\mathbf{x}}_{0})e^{-\alpha i}$ almost surely for $i$ large enough. Again by Lemma 3.4 we have that
@@
This shows that
$\sum_{i=0}^{\infty}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|\leq\frac{1}{\sqrt{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\sum_{i=0}^{\infty}\sqrt{f({\mathbf{x}}_{i})}=\sqrt{\frac{f({\mathbf{x}}_{0})}{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\frac{1}{1-e^{\alpha/2}}<\infty,$
and so the total length of the trajectory is finite. Thus ${\mathbf{x}}_{i}$ almost surely converges absolutely to a local minimum of $f$. ∎
@@ 4 Quantitative Results in Low Dimensions
The goal of this section is to study the hitting time $\tau$ of the sublevel set of the critical value. By construction, the stochastic process $x^{(i)}$ is a Markov chain, since the step $x^{(i+1)}-x^{(i)}$ depends only on the value of $x^{(i)}$. As such, for small enough step size $\eta$ we can approximate the algorithm by an appropriate Ito stochastic differential equation (SDE).
@@ 4.1 Two-Dimensional Euclidean Case
In the simplest case we consider the Morse function $f(x,y)=ax^{2}-by^{2}$ on $\mathbb{R}^{2}$ in the Euclidean metric. Due to the scale invariance we can normalize the state after each iteration and so we obtain a discrete-time stochastic process on the unit circle. For small enough stepsize $0<\eta\ll 1$ the iteration rule is approximately [^5] The formula is obtained by locally approximating the unit circle by its tangent line. given by
$\Delta\phi^{i}=\phi^{i+1}-\phi^{i}=\eta\,\frac{a+b}{2}\sin(2\phi)+u_{2}\eta\sqrt{(a\cos(\phi))^{2}+(b\sin(\phi))^{2}}$ (9)
where $\mathbf{u}=(u_{1},u_{2})$ is chosen uniformly at random from the unit circle and hence
$\mathbb{E}[\Delta\phi]=\eta\,\frac{a+b}{2}\sin(2\phi),\quad\operatorname{Var}(\Delta\phi)=\frac{\eta\sqrt{(a\cos(\phi))^{2}+(b\sin(\phi))^{2}}}{2}$
Using [32, Sec. 6.2] we can approximate this process using the stochastic differential equation (SDE)
$d\Phi_{t}=\frac{a+b}{2}\sin(2\phi)dt+\frac{\sqrt{\eta((a\cos(\phi))^{2}+(b\sin(\phi))^{2})}}{2}dW_{t}$ (10)
For simplicity we consider the simplest case $a=b=1$. Then the limiting SDE reads
$d\Phi_{t}=\sin(2\Phi_{t})dt+\frac{\sqrt{\eta}}{2}dW_{t}$
The approximation is visualized in Figure 3.
Figure 3: We use the constants $\eta=0.01$, $T=10$, $a=b=1$. Left: $50$ realizations of the stochastic process (9) for $T/\eta=1000$ steps. Right: $50$ realizations of the stochastic process (10) for time $T$ and with timestep size $\Delta t=0.001$ (simulated using the Euler-Maruyama scheme [32, Sec. 9.1]).
In order to understand how long it takes the algorithm to pass the saddle point we are interested computing in the hitting time
$\tau=\inf\{t>0:\Phi_{t}=\pm\frac{\pi}{4}\}.$
Close to the origin, the SDE can be linearized as $d\Phi_{t}=2\Phi_{t}dt+\frac{\sqrt{\eta}}{2}dW_{t}$, which is a mean repelling Ornstein–Uhlenbeck process. Away from the origin, the deterministic part dominates and we can solve the (deterministic) ODE $\dot{\Phi}_{t}=\sin(2\Phi_{t})$. First we approximate the hitting time distribution of the repelling Ornstein-Uhlenbeck process.
@@
Let $X_{t}$ be the solution of the SDE $dX_{t}=\kappa X_{t}dt+\sigma dW_{t}$ with $\kappa,\sigma>0$ and $X_{0}=0$. Setting $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, it holds that $X_{t}\sim\mathcal{N}(0,\tilde{\sigma}(t)^{2})$. If we denote by $\tau_{c}$ the hitting time of ±$\pm c$ (where $c>0$), we find the lower bound $\Pr[\tau_{c}\leq t]\geq\Pr[|X_{t}|\geq c]=1+\operatorname{erf}\big{(}\frac{-c}{\tilde{\sigma}(t)\sqrt{2}}\big{)}$ where $\operatorname{erf}$ denotes the error function.
@@
where $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, see [32, Sec. 4.2 & 4.4]. It is clear that if $|X_{t}|\geq c$ then $\tau_{c}\leq t$, which implies the last statement. ∎
The approximation is very accurate with error smaller than $\Delta t$ if $\kappa\Delta t\geq 1$ and $c\gg\sigma/\sqrt{\kappa}$, as illustrated in Figure 4.
As in the proof of Lemma 4.1 the solution of the SDE with initial condition $X_{0}=c$ is given by
@@
Hence for the error in $\tau_{c}$ to be small compared to $\Delta t$, we need $c\gg\frac{\sigma}{\sqrt{\kappa}}$ if $\kappa\Delta t$ is big, and $c\gg\frac{\sigma}{\kappa\sqrt{\Delta t}}$ if $\kappa\Delta t$ is small.
Figure 4: We used the constants $\kappa=2$, $\sigma=3$ and $c=10$. Shown are the histograms of the c.d.f. and p.d.f of $\tau_{c}$ computed using timesteps of size $0.001$ (orange) and the lower bound for the c.d.f. and the resulting p.d.f. obtained in Lemma 4.1 (blue).
@@
Now if we choose $c$ small enough that the linearization of $\sin$ is accurate but large enough that the approximation of Lemma 4.1 is good (which is always possible if $\eta$ is small enough), then we can approximate $\tau$ as follows.
@@
We fix some value $c$ and note that the ODE $\dot{\phi}_{t}=\sin(2\phi_{t})$ with initial condition $\phi_{0}=c$ has the solution $\phi_{t}=\arctan(e^{2t}\tan(c))$, and hence the hitting time of $\pi/4$ is $\tilde{\tau}_{c}=-\tfrac{1}{2}\ln(\tan(c))$. Hence we have approximately $\tau\simeq\tau_{c}+\tilde{\tau}_{c}$. This is illustrated in Figure 5 for realistic values.
@@
Figure 5: Using $\eta=0.01$ we plot the c.d.f. of the hitting time $\tau$ of $\tfrac{\pi}{4}$. In orange for the difference equation, in blue for the stochastic differential equation ($\Delta t=0.001$), and in red for the analytic approximation.
5 Ground State Optimization

### v2

@@ 3.2 Escaping Saddle Points
Corollary 3.13.
Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function on $M$. For any strict saddle point ${\mathbf{x}}\in M$ with value $f({\mathbf{x}})=c$ there exists $N>0$ and a neighborhood $U$ of ${\mathbf{x}}$ such that
${\Pr(f({\mathbf{x}}_{i+N})<c\,|\,{\mathbf{x}}_{i},\ldots,{\mathbf{x}}_{i+N}\in U)}>\delta.$
Proof.
This follows immediately from Cor. 3.10, Lem. 3.11 and Lem. 3.12. ∎
@@
Let $M$ be a Riemannian manifold and let $f$ be a Morse–Bott function on $M$. Further assume that $f:M\to\mathbb{R}$ has compact sublevel sets and that $\mathbf{x}_{0}\in M$ is not a critical point of $f$. Then the probability that $\mathbf{x}_{i}$ converges to the set of strict saddle points as $i\to\infty$ is zero.
@@
Let $C$ be the set of all strict saddle points $\mathbf{x}$ with $f(\mathbf{x})\leq f(\mathbf{x}_{0})$. By assumption on $f$ the set $C$ is compact; actually, $C$ consists of finitely many connected compact submanifolds, cf. [57, Def. 2.41]. We will show that the probability that $\mathbf{x}_{i}$ converges to $C$ is zero. Note that by Lem. 3.7 the function value is (surely) not increasing, and therefore we can focus on $C_{c}:=C\cap f^{-1}$$(c)$ for some $c\leq f(\mathbf{x}_{0})$. By compactness of $C_{c}$ we can choose for each $\mathbf{z}\in C_{c}$ balls $B_{r_{z}}(\mathbf{z})$, $B_{R_{z}}(\mathbf{z})$ around $\mathbf{z}$ with $r_{z}<R_{z}$ such that both balls are contained in a neighborhood $U(\mathbf{z})$ of $\mathbf{z}$ as in Cor. 3.13 and such that any realization $\mathbf{x}_{i}$ needs at least $N$ steps to transverse the spherical shells $B_{R_{z}}(\mathbf{z})\setminus B_{r_{z}}(\mathbf{z})$. Since $C_{c}$ is compact we can find a finite covering by balls of the form $B_{r_{z}/2}(\mathbf{z})$, i.e.
$C_{c}\subset\bigcup_{k=1}^{K}B_{r_{k}/2}(\mathbf{z}_{k})\,,$
Next define $r_{*}:=\min_{k=1,\dots K}r_{k}/2$. Then, by assumption, there exists some $L\in\mathbb{N}$ such that $d(\mathbf{x}_{i},C_{c})<r_{*}$ for all $i\geq L$. Hence we conclude $d(\mathbf{x}_{L},\mathbf{z})\leq r_{*}$ for some $\mathbf{z}\in C_{c}$ and moreover $\mathbf{z}\in B_{r_{k}/2}(\mathbf{z}_{k})$ for some $k\in\{1,\dots,K\}$. Thus, the triangle inequality implies $\mathbf{x}_{L}\in B_{r_{k}}(\mathbf{z}_{k})$ and therefore we conclude $\mathbf{x}_{L},\mathbf{x}_{L+1},\dots,\mathbf{x}_{L+N}$ is in $U(\mathbf{z}_{k})$. Since $\mathbf{x}_{L+N}$ again satisfies $d(\mathbf{x}_{L+N},C_{c})<r_{*}$ we can repeat the above arguments and obtain $\mathbf{x}_{L+N},\mathbf{x}_{L+N+1},\dots,\mathbf{x}_{L+2N}$ are in some $U(\mathbf{z}_{k^{\prime}})$, etc. Combining this observation with Cor. 3.13 we get
$\Pr(\mathbf{x}_{i}\to C_{c}\;\text{for}\;i\to\infty)\leq\sum_{L=1}^{\infty}{\Pr}_{L}=0\,,$
where $\Pr_{L}$ is defined by
[…]
Above we used the fact that by Markovianity the behavior of each sequence of $N$ consecutive steps behaves independently. Since there are only finitely many critical submanifolds contained in $C$ it follows that there are only finitely many $c\leq f(\mathbf{x}_{0})$ such that $C_{c}\neq\emptyset$ and thus we conclude that the probability that $\mathbf{x}_{i}$ converges to $C$ is also zero. ∎
3.3 Almost Sure Convergence
The results proven so far guarantee that the algorithm (cp. Fig. 1) converges almost surely to the set of local minima:
Proposition 3.15.
Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with compact sublevel sets. Assume that $\mathbf{x}_{0}$ is not a critical point. Then, for stepsize $\eta\leq 1/\ell$, the randomized gradient descent algorithm converges almost surely to the set of local minima.
Proof.
Since by Lem. 3.7 the function value cannot increase, and since by compactness of sublevel sets the function is lower bounded, the function value must (surely) converge to some value $f^{\star}$. By Cor. 3.8, $f^{\star}$ is almost surely a critical value. By Lem. 3.14 the algorithm almost surely does not converge towards the set of strict saddle points. Hence it has converge almost surely to the set of local minima. ∎
In order to prove convergence to a single local minimum, we first need the following technical lemma.
@@ 3.3 Almost Sure Convergence
Let $\mathcal{E}_{i}$ be probability distributions over the interval $[0,1]$ and let $E_{i}\sim\mathcal{E}_{i}$ for $i\in\mathbb{N}$. Let $\varepsilon>0$ be given. We assume that there is $q>0$ such that
$\Pr(E_{i}\geq\varepsilon\,|\,E_{1},\ldots,E_{i-1})\geq q$
for all $i\in\mathbb{N}$. Then for any $\alpha\in(0,q\ln\tfrac{1}{1-\varepsilon})$ it holds that almost surely
@@
First recall the following basic fact. Consider a sequence of i.i.d. biased coin tosses $X_{i}\in\{0,1\}$ with $\Pr(X_{i}=1)=q$. By the strong law of large numbers [35, Thm. 20.1] the average $\overline{X}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}X_{i}$ converges almost surely to the expectation $q$. Hence for any $p\in(0,q)$, there almost surely exists some $n_{0}$ large enough such that $\overline{X}_{n}\geq p$ for all $n>n_{0}$.
Let any value $n$, and index set $I=\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,n\}$ (ordered increasingly) be given. Then, using the law of total probability,
@@
Finally this allows us to prove the main result:
@@
Let $M$ be a Riemannian manifold and let $f:M\to\mathbb{R}$ be an $\ell$-smooth Morse–Bott function with compact sublevel sets. Further assume that ${\mathbf{x}}_{0}$ is not a critical point of $f$. Then, for step size $\eta\leq 1/\ell$, the randomly projected gradient descent algorithm converges almost surely to a local minimum.
@@
The idea of the proof is as follows: first we show that almost surely the distance to the set of local minima decreases exponentially. Then this implies that the size of each step decreases exponentially as well, and hence the algorithm converges absolutely to a local minimum.
Prop. 3.15 shows that ${\mathbf{x}}_{i}$ converges almost surely to some set of local minima denoted $C$. Let ${\mathbf{z}}\in C$ be such a local minimum, and let $H({\mathbf{z}})$ denote the Hessian at ${\mathbf{z}}$, and let $a_{\min}({\mathbf{z}})$ denote the smallest non-zero eigenvalue of $H({\mathbf{z}})$. Assume w.l.o.g. that $f(\mathbf{z})=0$. By compactness and continuity $a_{\min}({\mathbf{z}})$ has a minimum on $C$, which we denote $a_{\min}$. In a sufficiently small neighborhood of ${\mathbf{z}}$, working in a chart as provided by Lem. 3.12, we obtain
@@
Together with Lem. 3.7 and Cor. 3.4 we find
$f({\mathbf{x}}_{i+1})\leq\Big{(}1-((\mathbf{u}_{i})_{N})^{2}\delta 2a_{\min}\Big{)}f({\mathbf{x}}_{i})$
with $\delta=\eta(1-\tfrac{\ell\eta}{2})<\tfrac{1}{2a_{\min}}$. By Lem. 3.16, applied to $E_{i}=((\mathbf{u}_{i})_{N})^{2}\delta a_{\min}$, this shows that there is some $\alpha>0$ such that $f({\mathbf{x}}_{i})\leq f({\mathbf{x}}_{0})e^{-\alpha i}$ almost surely for $i$ large enough. Again by Lem. 3.7 we have that
@@
This shows that for $L\in\mathbb{N}$ large enough we can assume that all $\mathbf{x}_{i}$ with $i\geq L$ are contained in a single chart of the same kind as in Lem. 3.7 and thus we conclude
$\sum_{i=L}^{\infty}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|\leq\frac{1}{\sqrt{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\sum_{i=L}^{\infty}\sqrt{f({\mathbf{x}}_{i})}\leq\sqrt{\frac{f({\mathbf{x}}_{0})}{\tfrac{1}{\eta}-\tfrac{\ell}{2}}}\frac{e^{L\alpha/2}}{1-e^{\alpha/2}}<\infty,$
and so the total length of the trajectory is finite. Thus ${\mathbf{x}}_{i}$ almost surely converges to a local minimum of $f$. ∎
@@ 4 Quantitative Results in Low Dimensions
The goal of this section is to study how long it takes to pass a saddle point. More precisely, we want to understand the hitting time $\tau$ of the sublevel set of the critical value at a saddle point. By construction, the discrete stochastic process ${\mathbf{x}}_{i}$ is a Markov chain, since the step ${\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}$ depends only on the value of ${\mathbf{x}}_{i}$, and not on previous steps. As such, for small enough step size $\eta$, we can approximate the algorithm by an appropriate Ito stochastic differential equation (SDE). As we will see, the accuracy of this approximation improves with smaller step size $\eta$, at the cost of slower convergence.
@@ 4.1 Two-Dimensional Euclidean Case
[…]
where $\mathbf{u}=(u_{1},u_{2})$ is chosen uniformly at random from the unit circle and hence using Lem. A.1 we find that
$\mathbb{E}[\Delta\phi_{i}]=\eta\,\sin(2\phi_{i}),\quad\operatorname{Var}(\Delta\phi_{i})=\frac{\eta}{2}\,.$
Using [38, Sec. 6.2] we can approximate this process using the stochastic differential equation (SDE)
$d\Phi_{t}=\sin(2\Phi_{t})dt+\frac{\sqrt{\eta}}{2}dW_{t}\,,$ (10)
where $W_{t}$ denotes a standard Brownian motion. The approximation is visualized in Fig. 2.
Figure 2: We use the step size $\eta=0.01$, total time $T=10$, and initial state $\phi_{0}=\Phi_{0}=0$. Left: $50$ realizations of the discrete-time stochastic process (9) for $T/\eta=1000$ steps. Right: $50$ realizations of the continuous-time stochastic process (10) for time $T=10$, simulated using the Euler–Maruyama scheme [38, Sec. 9.1] with time step size $\Delta t=0.001$.
In order to understand how long it takes the algorithm to pass the saddle point, we are interested in computing the hitting time
$\tau=\inf\{t\geq 0:\Phi_{t}=\pm\tfrac{\pi}{4}\},$
where we always assume that $\Phi_{0}=0$. Close to 00, the SDE can be linearized as $d\Phi_{t}=2\Phi_{t}dt+\tfrac{\sqrt{\eta}}{2}dW_{t}$, which is a mean-repelling Ornstein–Uhlenbeck process. Away from 00, the deterministic part dominates and we can solve the (deterministic) ODE $\dot{\Phi}_{t}=\sin(2\Phi_{t})$.
First we approximate the hitting time distribution of the mean-repelling Ornstein–Uhlenbeck process. More precisely, we find a lower bound on the cumulative distribution function (c.d.f.) of $\tau_{c}$, the hitting time of ±$\pm c$.
@@
Let $X_{t}$ be the solution of the SDE $dX_{t}=\kappa X_{t}dt+\sigma dW_{t}$ with $\kappa,\sigma>0$ and $X_{0}=0$. Setting $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, it holds that $X_{t}\sim\mathcal{N}(0,\tilde{\sigma}(t)^{2})$. If we denote by $\tau_{c}$ the hitting time of ±$\pm c$ (where $c>0$), we find the lower bound $\Pr[\tau_{c}\leq t]\geq\Pr[|X_{t}|\geq c]=1+\operatorname{erf}\big{(}\frac{-c}{\tilde{\sigma}(t)\sqrt{2}}\big{)}$ where $\operatorname{erf}(\cdot)$ denotes the error function.
@@
where $\tilde{\sigma}(t)=\sigma\sqrt{\frac{e^{2\kappa t}-1}{2\kappa}}$, see [38, Sec. 4.2 & 4.4]. It is clear that if $|X_{t}|\geq c$ then $\tau_{c}\leq t$, which implies the last statement. ∎
The approximation is very accurate, as illustrated in Fig. 3. The following remark gives some details on the quality of the approximation.
Figure 3: Lower bound for the hitting time $\tau_{c}$ of the mean-repelling Ornstein–Uhlenbeck process, as obtained in Lem. 4.1. We used the constants $\kappa=2$, $\sigma=3$ and $c=10$. Shown are the histograms of the cumulative distribution function (c.d.f.) and probability density function (p.d.f.) of $\tau_{c}$ for the SDE $dX_{t}=\kappa X_{t}dt+\sigma dW_{t}$ computed using time steps of size $0.001$ (orange), and the analytical lower bound for the c.d.f. and the resulting p.d.f. obtained in Lem. 4.1 (blue).
Remark 4.2.
As in the proof of Lem. 4.1 the solution of the SDE with initial condition $X_{0}=c$ is given by
@@
Hence, for the error in $\tau_{c}$ to be small compared to some given $t>0$, we need $c\gg\frac{\sigma}{\sqrt{\kappa}}$ if $\kappa t$ is big, and $c\gg\frac{\sigma}{\kappa\sqrt{t}}$ if $\kappa t$ is small.
@@
Now if we choose $c$ small enough that the linearization of $\sin(\cdot)$ is accurate, but large enough that the approximation of Lem. 4.1 is good (which is always possible if $\eta$ is small enough), then we can approximate $\tau$ as follows.
@@
We fix some value $c$ and note that the ODE $\dot{\phi}(t)=\sin(2\phi(t))$ with initial condition $\phi(0)=c$ has the solution $\phi(t)=\arctan(e^{2t}\tan(c))$, and hence the hitting time of $\pi/4$ is $\tilde{\tau}_{c}=-\tfrac{1}{2}\ln(\tan(c))$. Hence we have approximately $\tau\simeq\tau_{c}+\tilde{\tau}_{c}$. This is illustrated in Fig. 4 for realistic values.
@@
Figure 4: Using $\eta=0.05$ we plot the c.d.f. of the hitting time $\tau$ of $\tfrac{\pi}{4}$. In orange for the difference equation (9), in blue for the stochastic differential equation (10) (with time step $\Delta t=0.001$), and in red for the analytic approximation as described above.
5 Major Application:
Ground-State Optimization in Quantum Physics
Among the Riemannian optimization techniques adopted in quantum information science, hybrid quantum-classical algorithms [10] such as variational quantum algorithms (VQAs) [18, 69], have been developed to solve ground-state (smallest eigenvalue) problems [58].
From the perspective of quantum information and computation, in VQAs, a fixed parameterized quantum circuit, i.e. a parameterized set of unitary transformations (usually implemented by a fixed set of unitary gates), is iteratively optimized in tandem with a classical computer to minimize a cost function whose global optima encode the solution(s) to the desired problem. More precisely, a quantum computer, including a measurement device, is used to output the current (e.g., expectation) value of the cost function in each iterative step. Interleaved, a classical computer component provides the optimization routine that determines the parameter update for the next iteration on the quantum computer. Since quantum computers can, in principle, handle some classically exponentially complex problems, the hope is that VQAs will be able to solve problems of large size that a solely classical algorithm would be incapable of.
However, since the optimization problems VQAs aim to solve are highly non-linear, which can often be traced back to the quantum circuit parametrization [48, 43], the iterative search for the optimal parameters can get stuck in suboptimal solutions [12]. To overcome this challenge, adaptive quantum algorithms have been designed. Instead of fixing a quantum circuit and picking a parametrization as in VQAs, adaptive algorithms successively grow a quantum circuit based on measurement data from the quantum computer. In this approach, the cost function is minimized while the quantum circuit is grown, see [29, 73, 49, 50]. One of the most prominent examples of such a strategy is the so-called adapt-vqe algorithm [29], which was originally proposed to solve ground state problems in chemistry, and was further developed later to solve combinatorial optimization problems on quantum computers [68].
Unfortunately, since typically in adaptive quantum algorithms the circuit growth is informed by gradient estimates performed by the quantum computer, adaptive quantum algorithms face similar challenges as VQAs, i.e. the adaptive circuit growth can get stuck when gradients (asymptotically) vanish. This issue has recently been addressed by identifying adaptive quantum algorithms as quantum-computer implementations of Riemannian gradient flows on the special unitary group $\text{SU}(d)$ of dimension $d=2^{n}$ where $n$ is the number of qubits of the quantum computer [73]. Since implementing Riemannian gradient methods on quantum computers requires in general exponential resources (i.e. the number of gates or the number of iterations of the corresponding quantum circuit grows exponentially in $n$), [73] proposed projecting the Riemannian gradient into smaller dimensional subspaces that scale polynomially in $n$, which in turn yields scalable quantum-computer implementations. As already mentioned, “dimension reduction” for efficient quantum-computer implementations comes at the cost of convergence not only to local minima of the cost function but also to artificial minima induced by projections of the full Riemannian gradient into fixed subspaces. These artifacts can be avoided by “random projections” of the Riemannian gradient [50]. The random projections can be implemented on quantum computers as depicted in Figs. 1 and 5. More precisely, an efficient quantum-computer implementation of these random directions can be achieved through exact and approximate unitary 2-designs [25].
Figure 5: Schematic implementation of the discretized Riemannian gradient scheme by a quantum algorithm in case the manifold is $M=\mathrm{SU}(2^{n})$. Thus the dimension of the problem scales exponentially with the number $n$ of quantum particles (qubits). Each quantum circuit $U_{k}\in\mathrm{SU}(2^{n})$ is adaptively turned into $U_{k+1}=\exp\left(\frac{-i\eta\,dJ_{k}(0)}{d\theta}H_{k}\right)U_{k}$ by moving in a random tangent-space direction $iH_{k}$ whose corresponding projection $\frac{dJ_{k}(0)}{d\theta}=\langle\text{grad}J(U_{k}),iH_{k}U_{k}\rangle$ is measured on a quantum device. Here $\eta$ denotes the step size and $J_{k}(\theta)$ the cost function $\theta\mapsto J_{k}(\theta):=\mathrm{tr}(A\exp(-i\theta H_{k})U_{k}\rho U_{k}^{*}\exp(i\theta H_{k}))$ of the $k$-th iteration. The measurement of $J_{k}(\theta)$ is done on the quantum computer, and the statistics of repeated measurements to obtain the expectation value is done on a classical computer. In this way a randomized gradient flow as in Fig. 1 is implemented. In this work we show that for sufficiently small $\eta$ such a (discretized) flow converges almost surely to the global optimum.